# FlowerTune: A Cross-Domain Benchmark for Federated Fine-Tuning of Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) [1, 2, 7, 23, 55, 63] have exhibited remarkable performance across a broad spectrum of machine learning tasks and domains, including general natural language processing (NLP), medical question answering [50, 52], financial sentiment analysis [64], and code generation [29, 30]. These capabilities are typically attained through supervised fine-tuning applied to large-scale pre-trained models [20, 25, 29].
@@
Despite their success, the current paradigm of LLM development heavily depends on large volumes of publicly available data [1, 23, 55]. While it is widely acknowledged that increased data volume typically enhances model performance, recent studies have raised concerns that the supply of high-quality public data may be exhausted within a few years [53, 56]. Moreover, there is an increasing demand for specialized LLMs that can integrate domain-specific knowledge not readily accessible in publicly available web-based corpora—particularly in sensitive domains such as healthcare and finance. However, accessing data stored within relevant institutions and organizations raises substantial privacy and security concerns, and is further complicated by challenges related to data storage and transfer, and communication infrastructure.
@@
LLM fine-tuning via federated learning (FL) offers a promising solution to address these challenges [35, 66, 67]. FL enables collaborative model optimization by facilitating access to a wider range of sensitive datasets without the need to share raw data. This approach holds significant potential for the development of more fairness-preserving [40, 51] and domain-adapted LLMs. Recent studies [35, 66, 67] have begun to explore FL algorithms and fine-tuning strategies on selected models within this context. With rapid advancements and increasing accessibility of LLM pre-training, a broad array of pre-trained models are now available for downstream fine-tuning. However, their suitability for deployment in FL environments remains largely unexplored. Several critical questions persist: How effectively do these models perform under federated settings using existing aggregation and fine-tuning strategies? Are they compatible with the prevalent resource constraints in FL scenarios, such as communication overhead and memory limitations? And can FL offer a practical pathway for training more domain-specialized LLMs?
@@
To systematically address these open challenges and deepen understanding, we introduce the FlowerTune LLM Leaderboard  [^1] https://flower.ai/benchmarks/llm-leaderboard, a benchmarking initiative built on the Flower Platform [12], designed to evaluate the performance of various pre-trained LLMs under various federated fine-tuning scenarios across multiple domains. The leaderboard focuses on four high-impact application areas involving sensitive or private data: general NLP, finance, medical, and coding. Each domain includes a dedicated federated dataset for instruction tuning, along with domain-specific evaluation metrics. By establishing baseline results across a range of models and domains in an open-source, community-driven framework, the FlowerTune LLM Leaderboard has attracted an increasing number of submissions from both academic and industry contributors. Building upon these contributions, we conduct a first-of-its-kind comprehensive set of fine-tuning experiments to benchmark 26 base models under a unified FL setting. This enables valuable insights into the feasibility and effectiveness of deploying federated fine-tuning for LLMs, and would further assist in accelerating the development of more inclusive, privacy-preserving, and domain-specialized language models for real-world applications.
@@ 2.2 Federated fine-tuning
To emulate data distributions that could be expected across institutions such as medical, financial, and educational organizations, we carefully select four domain-specific public datasets to support the respective challenges. Specifically, we use alpaca-gpt4  [^2] https://huggingface.co/datasets/flwrlabs/alpaca-gpt4 [47] for the general NLP domain, fingpt-sentiment-train  [^3] https://huggingface.co/datasets/flwrlabs/fingpt-sentiment-train [64] for finance, medical-flashcards  [^4] https://huggingface.co/datasets/flwrlabs/medical-meadow-medical-flashcards [24] for the medical domain, and code-alpaca-20k  [^5] https://huggingface.co/datasets/flwrlabs/code-alpaca-20k [15] for coding. Each dataset includes domain-specific instruction prompts paired with corresponding answers, designed to train an LLM to function as an assistant within its respective domain. We employ Flower Datasets [36] to partition each dataset into multiple shards of approximately equal size, simulating the data distribution across institutions in FL environments. A detailed statistical summary of the datasets and their corresponding splits is presented in Table 1, while representative example samples from each dataset are provided in the Appendix C.
@@
Challenges Fine-tuning dataset Total # samples # clients $len$(instruction) $len$(output)
General NLP alpaca-gpt4 [47] 52 K 20 60 677
Finance fingpt-sentiment-train [64] 76.8 K 50 112 9
Medical medical-flashcards [24] 34 K 20 92 349
Code code-alpaca-20k [15] 20 K 10 74 197
@@
General NLP MMLU [26] (STEM, Humanities, Social Sciences) MQA Accuracy
Finance FPB [42], FIQA [41], TFNS [70] Classification Accuracy
Medical PubMedQA [32], MedMCQA [45], MedQA [31], CareQA [8] MQA Accuracy
Code MBPP [9], HumanEval [16], MultiPL-E (JS, C++) [14] Code generation Pass@1
@@ 2.3 Downstream evaluation
For each domain, we provide a corresponding evaluation pipeline that includes domain-specific evaluation datasets (Table 2). Specifically, in the general NLP challenge, the MMLU dataset [26] is used for evaluation. MMLU consists of multiple-choice questions covering a broad range of subjects across STEM, humanities, and social sciences. The evaluation metric is accuracy, reflecting the performance of the tested LLM. For the finance challenge, evaluation is conducted using a sentiment classification pipeline on financial reports and social media posts. Three datasets are utilized: FPB [42], FIQA [41], and TFNS [70]. Model performance is assessed using accuracy as the evaluation metric. In the medical domain, we leverage four datasets—PubMedQA [32], MedMCQA [45], MedQA [31], and CareQA [8]—to evaluate the performance of the tuned LLM as a medical assistant. Each dataset consists of multiple-choice questions covering various medical topics. Evaluation is based on accuracy, reflecting the model’s ability to select the correct answers. For the coding challenge, we select MBPP [9], HumanEval [16], and MultiPL-E (JavaScript, C++) [14] as evaluation datasets. These datasets are specifically designed to assess code generation capabilities, and the pass@1 score is used as the evaluation metric, measuring the proportion of correctly generated programs on the first attempt.
@@ 3 Experimental settings
The FlowerTune LLM Leaderboard has attracted significant interest from the community, evidenced by a growing number of submissions. These submissions, together with the benchmarking code  [^6] https://github.com/yan-gao-GY/flowertune-benchmark and pipelines provided by FlowerTune, were instructive in allowing us to systematically conduct a first-of-its-kind comprehensive set of fine-tuning experiments to benchmark various base models within a unified FL setting. In addition to investigating different base models, we also explore the impact of diverse aggregation algorithms and fine-tuning strategies. Detailed descriptions of the experimental settings are provided in the following sections.
@@ 3.2 Federated learning settings
In the investigation of FL strategies, we select four popular algorithms: (1) FedAvg [43]: Federated Averaging (FedAvg) aggregates local model updates by computing their weighted average based on client data size. This method serves as a foundational method for many FL algorithms. (2) FedProx [37]: Federated Proximal (FedProx) extends FedAvg by introducing a proximal term to the local objective, which helps stabilize training when client data is non-IID by discouraging local models from drifting too far from the global model. (3) FedAvgM [27]: FedAvg with Momentum (FedAvgM) enhances FedAvg by incorporating server-side momentum during aggregation, which accelerates convergence and improves performance in heterogeneous environments. (4) FlexLoRA [10]: FlexLoRA is a communication-efficient FL method that integrates Low-Rank Adaptation (LoRA) into local training and employs adaptive aggregation to fuse updates, which improves scalability and personalization.
@@ 3.3 Local fine-tuning hyper-parameters
During the local fine-tuning phase, we adopt parameter-efficient fine-tuning (PEFT) techniques [20, 25], which substantially reduce computational requirements and accelerate the training process. Specifically, we investigate the following two popular PEFT methods in our federated LLM fine-tuning pipeline: (1) LoRA [28] introduces a PEFT method by injecting trainable low-rank matrices into pre-trained model weights, enabling a significant reduction in trainable parameters and computational cost without additional inference latency. (2) DoRA [38] builds upon LoRA by decomposing weights into magnitude and direction components, applying LoRA specifically to the directional part to better emulate the learning capacity of full fine-tuning while preserving efficiency.
In the local tuning process, the base model remains frozen, while only the LoRA/DoRA adapters are trained and communicated between clients and the central server. We employ DoRA while benchmarking various base models and additionally compare its performance against LoRA on selected models. Additionally, we integrate FlashAttention-2 [18] to further enhance training speed and computational efficiency. All experiments are conducted on NVIDIA A100 SXM4 (80 GB) GPUs. The complete set of training hyper-parameters is summarized in the Appendix A.2.
@@ 4.2 System performance analysis
VRAM consumption is influenced by several factors, including the architecture and size of the base model, the adapter size, and the input sequence length determined by the training dataset. Memory usage is often regarded as a primary bottleneck in federated on-device training. Experimental results demonstrate that the memory footprint of all evaluated models using FlowerTune tools remains below 80 GB—sufficient to fit within a single modern A100/H100 GPU, or across two mid-tier GPUs. In particular, larger models (e.g., Gemma2-9B-Instruct) demonstrate superior performance but incur higher memory consumption. Conversely, smaller models—such as those in the SmolLM2 family—achieve acceptable performance on specific challenges while requiring approximately 10 GB or less of VRAM. These trade-offs between performance and resource efficiency should be carefully considered when deploying models in real-world applications (see Appendix B.2).
@@
Table 8: Comparison of two fine-tuning techniques using the Qwen2.5-7B base model across the four challenges. FedAvg is used for aggregation. “SS” refers to Social Sciences. For the medical challenge, “PMQA”, “MMCQA”, “MQA” and “CQA” correspond to PubMedQA, MedMCQA, MedQA, and CareQA, respectively. For the coding challenge, “HE”, “M-JS” and “M-C++” represent HumanEval, MultiPL-E (JS), and MultiPL-E (C++). Communication (Comm.) and VRAM (Mem.) costs are reported in gigabytes.
@@
Methods STEM SS Humanities Avg Comm. Mem. PMQA MMCQA MQA CQA Avg Comm. Mem.
@@ 5 Related work
Instruction tuning is a form of fine-tuning that aligns LLMs with human intent using instruction-response datasets. Unlike task-specific fine-tuning, it promotes generalization across diverse tasks by training models to follow natural language prompts [71]. InstructGPT demonstrated that human-written instructions improve model helpfulness and safety [44], while Self-Instruct reduced data collection costs by using LLMs to generate synthetic instructions [58]. This approach enabled large-scale datasets like Super-NaturalInstructions and inspired methods such as Evo-Instruct and Instruct-SkillMix [33, 57, 69]. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA, address the high cost of fine-tuning by updating only small portions of model parameters [28]. This makes instruction tuning feasible in low-resource environments. Recent studies integrate LoRA into instruction workflows for improved scalability [11, 19], while advanced variants like ALoRA optimize performance by dynamically adjusting parameter allocation [39].
Federated LLM fine-tuning enables collaborative training across decentralized data sources. Recent frameworks such as OpenFedLLM [67], FederatedScope-LLM [35], and FATE-LLM [21] support instruction tuning, parameter-efficient methods, and industrial deployment. Benchmark FedLLM-Bench [66] and data enrichment tools such as FedIT-U2S [65] and FedDCA [59] provide datasets and pipelines for empirical studies and domain adaptation. Despite progress, federated LLM fine-tuning faces challenges including limited resources, system heterogeneity, and communication constraints. LoRA-based parameter-efficient fine-tuning has emerged as a key solution. Methods such as FedBiOT [61], FedALT [13], and FFA-LoRA [54] enhance efficiency by tuning lightweight adapters and reducing communication. Others—such as pFedLoRA [68], FLoRA [60], and FDLoRA [49]—support heterogeneous models and balance global vs. personalized knowledge. Privacy and adaptability are further addressed by FedRand [46] and LoRA-A2 [34]. Beyond adapter tuning, approaches like FedMKT [22] and FedBis [62] focus on personalization, knowledge transfer, and user preference alignment, reflecting increasing attention to robust and user-aware federated LLM development. However, existing work has not systematically examined the impact of diverse base models on federated fine-tuning, particularly in terms of downstream performance and system-level behavior—an essential but underexplored aspect of this emerging field.

### v2

@@ 1 Introduction
Large language models (LLMs) [abdin2024phi, abouelenin2025phi, allal2025smollm2, grattafiori2024llama, gemma_2024, yang2024qwen2] have exhibited remarkable performance across a broad spectrum of machine learning tasks and domains, including general natural language processing (NLP), medical question answering [saab2024capabilities, singhal2025toward], financial sentiment analysis [yang2023fingpt], and code generation [huang2023agentcoder, islam2024mapcoder]. These capabilities are typically attained through supervised fine-tuning applied to large-scale pre-trained models [ding2023parameter, han2024parameter, huang2023agentcoder].
@@
Despite their success, the current paradigm of LLM development heavily depends on large volumes of publicly available data [abdin2024phi, grattafiori2024llama, gemma_2024]. While it is widely acknowledged that increased data volume typically enhances model performance, recent studies have raised concerns that the supply of high-quality public data may be exhausted within a few years [spiceworks2024llmdata, villalobos2022will]. Moreover, there is an increasing demand for specialized LLMs that can integrate domain-specific knowledge not readily accessible in publicly available web-based corpora—particularly in sensitive domains such as healthcare and finance. However, accessing data stored within relevant institutions and organizations raises substantial privacy and security concerns, and is further complicated by challenges related to data storage and transfer, and communication infrastructure.
@@
LLM fine-tuning via federated learning (FL) offers a promising solution to address these challenges [kuang2024federatedscope, ye2024fedllm, ye2024openfedllm]. FL enables collaborative model optimization by facilitating access to a wider range of sensitive datasets without the need to share raw data. This approach holds significant potential for the development of more fairness-preserving [lyu2020collaborative, shi2023towards] and domain-adapted LLMs. Recent studies [kuang2024federatedscope, ye2024fedllm, ye2024openfedllm] have begun to explore FL algorithms and fine-tuning strategies on selected models within this context. With rapid advancements and increasing accessibility of LLM pre-training, a broad array of pre-trained models are now available for downstream fine-tuning. However, their suitability for deployment in FL environments remains largely unexplored. Several critical questions persist: How effectively do these models perform under federated settings using existing aggregation and fine-tuning strategies? Are they compatible with the prevalent resource constraints in FL scenarios, such as communication overhead and memory limitations? And can FL offer a practical pathway for training more domain-specialized LLMs?
@@
To systematically address these open challenges and deepen understanding, we introduce the FlowerTune LLM Leaderboard  [^1] https://flower.ai/benchmarks/llm-leaderboard, a benchmarking initiative built on the Flower Platform [beutel2020flower], designed to evaluate the performance of various pre-trained LLMs under various federated fine-tuning scenarios across multiple domains. The leaderboard focuses on four high-impact application areas involving sensitive or private data: general NLP, finance, medical, and coding. Each domain includes a dedicated federated dataset for instruction tuning, along with domain-specific evaluation metrics. By establishing baseline results across a range of models and domains in an open-source, community-driven framework, the FlowerTune LLM Leaderboard has attracted an increasing number of submissions from both academic and industry contributors. Building upon these contributions, we conduct a first-of-its-kind comprehensive set of fine-tuning experiments to benchmark 26 base models under a unified FL setting. This enables valuable insights into the feasibility and effectiveness of deploying federated fine-tuning for LLMs, and would further assist in accelerating the development of more inclusive, privacy-preserving, and domain-specialized language models for real-world applications.
@@ 2.2 Federated fine-tuning
To emulate data distributions that could be expected across institutions such as medical, financial, and educational organizations, we carefully select four domain-specific public datasets to support the respective challenges. Specifically, we use alpaca-gpt4  [^2] https://huggingface.co/datasets/flwrlabs/alpaca-gpt4 [peng2023instruction] for the general NLP domain, fingpt-sentiment-train  [^3] https://huggingface.co/datasets/flwrlabs/fingpt-sentiment-train [yang2023fingpt] for finance, medical-flashcards  [^4] https://huggingface.co/datasets/flwrlabs/medical-meadow-medical-flashcards [han2023medalpaca] for the medical domain, and code-alpaca-20k  [^5] https://huggingface.co/datasets/flwrlabs/code-alpaca-20k [codealpaca] for coding. Each dataset includes domain-specific instruction prompts paired with corresponding answers, designed to train an LLM to function as an assistant within its respective domain. We employ Flower Datasets [flwr_datasets] to partition each dataset into multiple shards of approximately equal size, simulating the data distribution across institutions in FL environments. A detailed statistical summary of the datasets and their corresponding splits is presented in Table 1, while representative example samples from each dataset are provided in the Appendix C. Additionally, a quantitative analysis of the client-level data distribution is provided in Appendix A.1. Although the number of samples is relatively balanced across clients within each domain, semantic similarity metrics reveal notable differences among clients, particularly in the general NLP, medical, and code domains.
@@
Challenges Fine-tuning dataset Total # samples # clients lenlen(instruction) lenlen(output)
General NLP alpaca-gpt4 [peng2023instruction] 52 K 20 60 677
Finance fingpt-sentiment-train [yang2023fingpt] 76.8 K 50 112 9
Medical medical-flashcards [han2023medalpaca] 34 K 20 92 349
Code code-alpaca-20k [codealpaca] 20 K 10 74 197
@@
General NLP MMLU [hendryckstest2021] (STEM, Humanities, Social Sciences) MQA Accuracy
Finance FPB [Malo2014GoodDO], FIQA [maia201818], TFNS [twitterfinancialsentiment] Classification Accuracy
Medical PubMedQA [jin2019pubmedqa], MedMCQA [pmlr-v174-pal22a], MedQA [jin2021disease], CareQA [arias-duart-etal-2025-automatic] MQA Accuracy
Code MBPP [austin2021program], HumanEval [chen2021evaluating], MultiPL-E (JS, C++) [cassano2023multipl] Code generation Pass@1
@@ 2.3 Downstream evaluation
For each domain, we provide a corresponding evaluation pipeline that includes domain-specific evaluation datasets (Table 2). Specifically, in the general NLP challenge, the MMLU dataset [hendryckstest2021] is used for evaluation. MMLU consists of multiple-choice questions covering a broad range of subjects across STEM, humanities, and social sciences. The evaluation metric is accuracy, reflecting the performance of the tested LLM. For the finance challenge, evaluation is conducted using a sentiment classification pipeline on financial reports and social media posts. Three datasets are utilized: FPB [Malo2014GoodDO], FIQA [maia201818], and TFNS [twitterfinancialsentiment]. Model performance is assessed using accuracy as the evaluation metric. In the medical domain, we leverage four datasets—PubMedQA [jin2019pubmedqa], MedMCQA [pmlr-v174-pal22a], MedQA [jin2021disease], and CareQA [arias-duart-etal-2025-automatic]—to evaluate the performance of the tuned LLM as a medical assistant. Each dataset consists of multiple-choice questions covering various medical topics. Evaluation is based on accuracy, reflecting the model’s ability to select the correct answers. For the coding challenge, we select MBPP [austin2021program], HumanEval [chen2021evaluating], and MultiPL-E (JavaScript, C++) [cassano2023multipl] as evaluation datasets. These datasets are specifically designed to assess code generation capabilities, and the pass@1 score is used as the evaluation metric, measuring the proportion of correctly generated programs on the first attempt.
@@ 3 Experimental settings
The FlowerTune LLM Leaderboard has attracted significant interest from the community, as evidenced by a growing number of submissions. These submissions, together with the benchmarking code  [^6] https://github.com/yan-gao-GY/flowertune-benchmark and pipelines provided by FlowerTune, were instructive in allowing us to systematically conduct a first-of-its-kind comprehensive set of fine-tuning experiments to benchmark various base models within a unified FL setting. In addition to investigating different base models, we also explore the impact of diverse aggregation algorithms and fine-tuning strategies. Detailed descriptions of the experimental settings are provided in the following sections.
@@ 3.2 Federated learning settings
In the investigation of FL strategies, we select four popular algorithms: (1) FedAvg [mcmahan2017communication]: Federated Averaging (FedAvg) aggregates local model updates by computing their weighted average based on client data size. This method serves as a foundational method for many FL algorithms. (2) FedProx [li2020federated]: Federated Proximal (FedProx) extends FedAvg by introducing a proximal term to the local objective, which helps stabilize training when client data is non-IID by discouraging local models from drifting too far from the global model. (3) FedAvgM [hsu2019measuring]: FedAvg with Momentum (FedAvgM) enhances FedAvg by incorporating server-side momentum during aggregation, which accelerates convergence and improves performance in heterogeneous environments. (4) FlexLoRA [bai2024federated]: FlexLoRA is a communication-efficient FL method that integrates Low-Rank Adaptation (LoRA) into local training and employs adaptive aggregation to fuse updates, which improves scalability and personalization.
@@ 3.3 Local fine-tuning hyper-parameters
During the local fine-tuning phase, we adopt parameter-efficient fine-tuning (PEFT) techniques [ding2023parameter, han2024parameter], which substantially reduce computational requirements and accelerate the training process. Specifically, we investigate the following two popular PEFT methods in our federated LLM fine-tuning pipeline: (1) LoRA [hu2022lora] introduces a PEFT method by injecting trainable low-rank matrices into pre-trained model weights, enabling a significant reduction in trainable parameters and computational cost without additional inference latency. (2) DoRA [liu2024dora] builds upon LoRA by decomposing weights into magnitude and direction components, applying LoRA specifically to the directional part to better emulate the learning capacity of full fine-tuning while preserving efficiency. Both methods are applied with quantization, resulting in a QLoRA-style training approach [dettmers2023qlora].
In the local tuning process, the base model remains frozen, while only the LoRA/DoRA adapters are trained and communicated between clients and the central server. We employ DoRA while benchmarking various base models and additionally compare its performance against LoRA on selected models. Additionally, we integrate FlashAttention-2 [dao2023flashattention] to further enhance training speed and computational efficiency. All experiments are conducted on NVIDIA A100 SXM4 (80 GB) GPUs. The complete set of training hyper-parameters is summarized in the Appendix A.3.
@@ 4.2 System performance analysis
VRAM consumption is influenced by several factors, including the architecture and size of the base model, the adapter size, and the input sequence length determined by the training dataset. Memory usage is often regarded as a primary bottleneck in federated on-device training. Experimental results demonstrate that the memory footprint of all evaluated models using FlowerTune tools remains below 80 GB—sufficient to fit within a single modern A100/H100 GPU, or across two mid-tier GPUs. In particular, larger models (e.g., Gemma2-9B-Instruct) demonstrate superior performance but incur higher memory consumption. Conversely, smaller models—such as those in the SmolLM2 family—achieve acceptable performance on specific challenges while requiring approximately 10 GB or less of VRAM. This enables the training of these smaller models on edge devices, such as an NVIDIA Jetson board, MacBook, or Raspberry Pi with 16 GB of RAM, when considering memory consumption alone. However, computational speed represents another critical factor, which can only be accurately assessed on actual edge devices and is generally considerably slower than GPU-based training. These trade-offs between performance and resource efficiency should be carefully considered when deploying models in real-world applications (see Appendix B.2).
@@
Table 8: Comparison of two fine-tuning techniques using the Qwen2.5-7B base model across four challenges. FedAvg is used for aggregation. “SS” refers to Social Sciences. For the medical challenge, “PMQA”, “MMCQA”, “MQA” and “CQA” correspond to PubMedQA, MedMCQA, MedQA, and CareQA, respectively. For the coding challenge, “HE”, “M-JS” and “M-C++” represent HumanEval, MultiPL-E (JS), and MultiPL-E (C++). Communication (Comm.) and VRAM (Mem.) costs are reported in gigabytes.
@@
Methods STEM $S$ Humanities Avg Comm. Mem. PMQA MMCQA MQA CQA Avg Comm. Mem.
@@ 5.1 LLM instruction tuning
Instruction tuning is a form of fine-tuning that aligns LLMs with human intent using instruction-response datasets. Unlike task-specific fine-tuning, it promotes generalization across diverse tasks by training models to follow natural language prompts [zhang2023instruction]. InstructGPT demonstrated that human-written instructions improve model helpfulness and safety [ouyang2022training], while Self-Instruct reduced data collection costs by using LLMs to generate synthetic instructions [wang2023self]. This approach enabled large-scale datasets like Super-NaturalInstructions and inspired methods such as Evo-Instruct and Instruct-SkillMix [kaurinstruct, wang2022super, zeng2024automatic]. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA, address the high cost of fine-tuning by updating only small portions of model parameters [hu2022lora]. This makes instruction tuning feasible in low-resource environments. Recent studies integrate LoRA into instruction workflows for improved scalability [balne2024parameter, dettmers2023qlora], while advanced variants like ALoRA optimize performance by dynamically adjusting parameter allocation [liu2024alora]. Beyond supervised tuning, reinforcement learning from human feedback (RLHF) aligns model behavior with human preferences, as seen in the training of ChatGPT [bai2022training]. While effective, RLHF faces challenges in reward modeling, feedback quality, and scalability [ziegler2019fine, lee2023rlaif].
Recent work has explored efficient on-device fine-tuning of LLMs to address privacy and resource constraints. GSQ-Tuning [zhou2025gsq] enables fully integer-based on-device LLM fine-tuning using group-shared exponents, while QEFT [lee2024qeft] improves efficiency by updating only sensitive weight columns in mixed precision. These methods extend PEFT methods like LoRA [hu2022lora] and QLoRA [dettmers2023qlora] for resource-constrained settings, and are applicable to federated cross-device fine-tuning. While our study focuses on cross-silo fine-tuning, we acknowledge that recent advances in on-device fine-tuning offer promising directions that could further enhance privacy and efficiency in federated settings, and we leave their exploration to future work.
5.2 Federated LLM fine-tuning
Federated LLM fine-tuning facilitates privacy-preserving training across decentralized data sources. Several frameworks and benchmarks have been developed to support this paradigm. OpenFedLLM [ye2024openfedllm] introduces a general framework for instruction tuning and alignment of LLMs without sharing raw data. FederatedScope-LLM [kuang2024federatedscope] provides a modular package supporting PEFT and automated benchmarking. FATE-LLM [fan2023fate] focuses on industrial deployment, supporting both horizontal and vertical federated settings. FedLLM-Bench [ye2024fedllm] addresses the benchmarking gap by offering realistic datasets, fine-tuning methods, and evaluation pipelines to support empirical studies. To enrich training data, FedIT-U2S [ye2024leveraging] transforms unstructured client data into instruction-response pairs, while FedDCA [zhang2024feddca] augments domain coverage to improve model generalization.
Despite these advancements, federated LLM fine-tuning faces challenges such as resource constraints, system heterogeneity, personalization, and communication efficiency. Parameter-efficient fine-tuning methods, particularly LoRA-based techniques, have become central to addressing these issues. FedBiOT [wu2024fedbiot] addresses limited resources by enabling clients to tune lightweight adapters instead of full models. FedALT [bian2025fedalt] improves LoRA performance in federated contexts by stabilizing client updates via adaptive training. Further improvements in LoRA are seen in FFA-LoRA [sun2024improving], which freezes select LoRA matrices to reduce communication, and in pFedLoRA [yi2023pfedlora], which supports model heterogeneity using shared small adapters. FLoRA [wang2024flora] introduces aggregation strategies for heterogeneous LoRA modules, while FDLoRA [qi2024fdlora] deploys dual adapters for balancing global and personalized knowledge. FedRand [park2025fedrand] enhances privacy via randomized LoRA parameter updates, and LoRA-A2 [koo2024towards] adapts LoRA structure under extreme heterogeneity.
Beyond adapter tuning, several approaches focus on client-specific optimization. FedMKT [fan2025fedmkt] enhances generalization by transferring knowledge between small and large models. FedBis [wu2024client] integrates user preference alignment through binary selectors. These methods reflect growing attention to personalization and robustness in federated LLM fine-tuning. However, existing work has not systematically examined the impact of diverse base models on federated fine-tuning, particularly in terms of downstream performance and system-level behavior—an essential but underexplored aspect of this emerging field.